# A Survey of Text Watermarking in the Era of Large Language Models

### v3

@@ Abstract.
Text watermarking algorithms play a crucial role in the copyright protection of textual content, yet their capabilities and application scenarios have been limited historically. The recent developments in large language models (LLMs) have opened new opportunities for the advancement of text watermarking techniques. LLMs not only enhance the capabilities of text watermarking algorithms through their text understanding and generation abilities but also necessitate the use of text watermarking algorithms for their own copyright protection. This paper conducts a comprehensive survey of the current state of text watermarking technology, covering four main aspects: (1) an overview and comparison of different text watermarking techniques; (2) evaluation methods for text watermarking algorithms, including their success rates, impact on text quality, robustness, and unforgeability; (3) potential application scenarios for text watermarking technology; (4) current challenges and future directions for development. This survey aims to provide researchers with a thorough understanding of text watermarking technology, thereby promoting its further advancement.
Text Watermark, Large Language Models
$c$s: Computing methodologies Natural language processingccs: Security and privacy Social aspects of security and privacy
@@ 1. Introduction
Text watermarking involves embedding unique, imperceptible identifiers (watermarks) into textual content. These watermarks are designed to be robust yet inconspicuous, ensuring that the integrity and ownership of the content are preserved without affecting its readability or meaning. Historically, text watermarking has played a crucial role in various domains, from copyright protection and document authentication to preventing plagiarism and unauthorized content distribution (Kamaruddin et al., 2018). With the advancement of Large Language Models (LLMs), both the techniques and application scenarios of text watermarking have seen significant development. As shown in Figure 1(a), this primarily includes the construction of enhanced text watermarking algorithms using LLMs, the application of existing text watermarking algorithms to LLMs, and the exploration of text watermarking algorithms that are more closely integrated with LLMs. The flourishing development of LLMs has propelled a thriving research landscape within the realm of text watermarking, as depicted in Figure 1(b). Especially with the advent of ChatGPT, text watermarking has notably surged into a research fervor. To help better understanding the mutually beneficial relationship between LLMs and text watermarking, this paper provides a survey of text watermarking techniques in the era of large language models.
In the subsequent content of this section, we will separately discuss why text watermarking benefits the application of LLMs (section 1.1), why utilizing LLMs can lead to the development of superior text watermarking algorithms (section 1.2), and the contributions of this survey along with the organization of the following sections (section 1.3).
@@ 1.1. Why is Text Watermarking Beneficial for LLMs?
However, the extensive use of LLMs has also introduced a set of problems and challenges. Firstly, the rapid generation of high-quality text by LLMs can facilitate the rapid spread of false information (Megías et al., 2021). Secondly, the issue of intellectual property related to large models is of vital importance. This includes the copyright of datasets (Tang et al., 2023) used for training large models and the addition of intellectual property rights to prevent the extraction of knowledge from the models (Zhao et al., 2023b). If effective tagging and detection methods for LLM-generated text could be implemented, it would significantly aid in mitigating the aforementioned issues. Text watermarking emerges as a promising solution to address these challenges. By embedding a unique, identifiable, and non-obtrusive marker (watermark) within the LLM-generated text, watermarking can enable the tracking and attribution of content produced by LLMs.
@@ 1.2. Why are LLMs Beneficial for Text Watermarking?
One of the main challenges in text watermarking is embedding watermarks without altering the original meaning or readability of the text. This requirement presents a significant challenge, as traditional text watermarking methods often struggle to modify the text without changing its semantics (Atallah et al., 2001; Topkara et al., 2006a; Meral et al., 2009). This is due to the need for text watermarking algorithms to have a strong understanding and control over the text’s semantics. Here, Large Language Models emerge as a game-changer. With their advanced understanding of language semantics and context, LLMs enable more sophisticated text watermarking methods that seamlessly integrate watermarks with minimal compromise to the text’s original meaning (Abdelnabi and Fritz, 2021; Zhang et al., 2023a). This synergy allows for the development of watermarking techniques that are not only more effective but also more subtle, ensuring that the text remains as intended while still carrying the necessary watermark features.
@@ 1.3. Why a Survey for Text Watermarking in the Era of LLMs?
Although text watermarking technology and large language models can effectively enhance each other, for instance, text generated by LLMs can be watermarked using text watermarking algorithms (Brassil et al., 1995; Por et al., 2012; Rizzo et al., 2016; Munyer and Zhong, 2023; Yang et al., 2022, 2023; Yoo et al., 2023b), or LLMs themselves can be utilized to embed watermarks in texts (Abdelnabi and Fritz, 2021; Zhang et al., 2023a). Additionally, watermark algorithms can be directly incorporated during the text generation process of LLMs (Kirchenbauer et al., 2023a; Zhao et al., 2023a; Liu et al., 2023c, b; Ren et al., 2023; Wu et al., 2023). However, up to date, no studies have attempted to comprehensively explore and understand text watermarking from a broader perspective. The current surveys on text watermarking mainly focus on techniques prior to the era of large language models (Alkawaz et al., 2016; Kamaruddin et al., 2018).
Therefore, in this work, we provide the first comprehensive survey of text watermarking algorithms in the era of large language models, covering the detailed definition of text watermarking algorithms and the interconnections between different kinds of text watermarking methods. Given the complexity and diversity of text watermarking technology, we have also detailed how to evaluate text watermarking algorithms from different perspectives, including success rate, robustness, impact on text quality and unforgeability. Additionally, we have introduced the current application scenarios of text watermarking algorithms including copyright protection, fake news detection, and academic integrity. This survey can provide researchers with a high-level understanding of text watermarking algorithms and a comparison of the similarities and differences between different text watermarking algorithms. Researchers who merely wish to employ text watermarking technology can select the appropriate algorithms and application scenarios based on the introduction provided in this survey.
Organization of this survey. The remainder of this survey is organized as follows: Section 2 introduces the definition of text watermarking and the essential properties of text watermarking algorithms. Section 3 and Section 4 discuss two significant types of text watermarking algorithms: text watermarking for existing text and text watermarking for LLMs. Section 5 elaborates on the different evaluation perspectives of text watermarking algorithms, including success rate, impact on text quality, robustness, and unforgeability. Section 6 presents the application scenarios of text watermarking algorithms, covering copyright protection, academic integrity and fake news detection. Section 7 analyzes the challenges still faced by current text watermarking research and explores future research directions. Finally, Section 8 concludes the survey.
@@ 2.1. Text Watermarking Algorithms
The watermarked text, denoted as $\mathbf{t}$, can be derived in two ways: it may be a modified version of the input text $\mathbf{x}$, where $\mathbf{x}$ is the original text, or it can be a new text generated in response to $\mathbf{x}$, such as when $\mathbf{x}$ serves as a prompt for a Large Language Model. The watermark message, denoted as $w$, can be a zero-bit watermark, signifying merely its presence or absence, or a multi-bit watermark, embedding detailed, customized information. The term ‘watermark payload’ will be used henceforth to describe the quantity of information conveyed by the watermark message.
@@ 2.2. Key Characteristics of Text Watermarking Algorithms
To facilitate a unified understanding of the objectives in designing text watermarking algorithms, this section introduces the key characteristics that a watermarking algorithm should possess. These primarily include a high success rate of watermark detection, minimal impact on text quality, robustness of the detector against text modifications, and unforgeability.
@@
Low impact on text quality. We use $\mathcal{A}(\mathbf{x},None)$ to denote the text generated without adding a watermark. If $\mathbf{x}$ is the text to be modified, the output will be $\mathbf{x}$ itself. If $\mathbf{x}$ is a prompt for LLM, it would be the text generated by LLM without a watermark. A watermark algorithm that does not significantly affect text quality will the following condition:
@@
where $\mathcal{R}$ is a function evaluating text quality from multiple perspectives, as will be discussed in seciton 4. $\delta$ represents a threshold. If the difference in the evaluated scores of two texts is less than this threshold, they are considered to be of similar quality.
Robustness to watermark removal attack. We use an operation $\mathcal{U}$ to denote the watermark removal operations, which will be detailed in seciton 4. If a watermarking algorithm is robust against watermark removal attacks, it should satisfy the following conditions:
@@
Unforgeability. Unforgeability refers to the difficulty for a third party to counterfeit text watermarks. Typically, if the watermark’s generator $\mathcal{A}$ is acquired by an attacker, the watermark can certainly be forged. Therefore, the detailed definition of unforgeability is whether an attacker, unable to access the watermark’s generator, can counterfeit the watermark. This usually divides into two scenarios: in one, the attacker cannot obtain the detector $\mathcal{D}$ and is limited to watermark detection, known as the private detection scenario. In the second scenario, the attacker has acquired the watermark’s detector, referred to as the public detection scenario.
@@ 2.3. Taxonomy of Text Watermarking Algorithms
As illustrated in Figure 2, text watermarking algorithms can be broadly classified into two categories. The first category, Watermarking for Existing Text, focuses on embedding watermarks into pre-existing texts. Detailed in Section 3, this method typically employs semantically invariant transformations to incorporate watermarks seamlessly into the existing text.
The second category, Watermarking for Large Language Models (LLMs), involves alterations to the large language models. As we’ll explore in Section 4, this approach either introduces specific features into the training dataset or modifies the text generation process of LLMs. In essence, it creates watermarked text $\mathbf{t}$ in response to an input prompt $\mathbf{x}$.
@@ 3. Watermarking for Existing Text
Figure 3. A more illustrative explanation of various text watermarking methods. Watermarking for Existing Text (section 3) involves modifying existing text to embed watermarks, primarily through format-based approaches (section 3.1) such as white-space substitution, lexical-based approaches (section 3.2) such as synonym substitution, syntactic-based approaches (section 3.3) e.g. passivization, and generation-based approaches (section 3.4) that directly generated watermarked text through pretrained language models. Watermarking for LLMs (section 4) refers to embedding watermarks in Large Language Models, ensuring the text generated includes these watermarks, which can be implemented during training time (section 4.1), during logits generation (section 4.2), or during token sampling (section 4.3).
@@ 3.1. Format-based Watermarking
Format-based watermarking approaches are inspired by image watermarking technology (Begum and Uddin, 2020). It does not modify the content of the text but introduces changes to its format that are difficult for humans to detect, thereby embedding a watermark. For example, Brassil et al. (1995) proposed line-shift coding and word shift-coding techniques, achieved by vertically shifting the positions of text lines or horizontally shifting the locations of words within text lines. Correspondingly, the watermark detection process involves measuring the distance between adjacent text line profiles or between adjacent word column profiles to detect shifts. However, this approach is limited to embedding watermarks in image-formatted text and cannot truly return a text string with an embedded watermark. Considering this, various watermarking methods that rely on the insertion or replacement of Unicode codepoints have been introduced. Por et al. (2012) proposed a watermarking scheme named UniSpach, which inserts Unicode space characters into inter-sentence, inter-word, end-of-line and inter-paragraph spacings. A study by Rizzo et al. (2016) presented a unicode homoglyph substitution text watermarking method. It exploits the fact that text symbols that are similar in appearance could have different Unicode codepoints. For instance, both U+0043 and U+216d visually represent letter ‘C’, while U+004c and U+216c appear as letter ‘L’. Following this, a family of simple watermarks named EASYMARK (Sato et al., 2023) is proposed recently, composed of three different methods: WHITEMARK, VARIANTMARK and PRINTMARK. Specifically, WHITEMARK replaces a whitespace (U+0020) with another codepoint of a whitespace (e.g. U+2004). VARIANTMARK leverages variation selectors of Unicode to embed hard-to-perceive format into CJK texts. PRINTMARK, coping with printed texts, uses ligature or whitespaces with slightly different lengths to embed watermark messages. Correspondingly, the watermark detection process involves searching for and counting the certain codepoints that have been inserted within the text. As these watermarking methods relied on the richness of Unicode encoding, their watermark payload are often quite large.
However, although these format-based watermarking methods allow for the simple and effective embedding of large payload watermarks in text without altering its specific content, modifications to the format can be easily spotted in certain scenarios, as mentioned by Por et al. (2012) in the case of DASH attack. As a result, these specially designed formats could be effortlessly removed through canonicalization (Boucher et al., 2022), such as reseting line spacing, searching and replacing certain codepoints through the entire text, etc. The spotted formats might also be used to forge watermark, further leading to the failure of effective detection.
@@ 3.2. Lexical-based Watermarking
Format-based approaches only modify the surface format of the text, making them easily spotted and, consequently, more vulnerable to targeted removal by reformatting. Therefore, it becomes imperative to explore alternative methods that enable deeper insertion of watermarks within text. Several studies applied word-level modifications by replacing selected words with their alternatives without changing the sentence syntax structure (Topkara et al., 2006b; Fellbaum, 1998; Munyer and Zhong, 2023; Yang et al., 2022, 2023; Yoo et al., 2023b). We refer to these methods as lexical-based watermarking approaches. Topkara et al. (2006b) presented a synonym substitution text watermarking approach, using the linguistic database WordNet(Fellbaum, 1998) as its synonym dictionary. The watermark detection process essentially replicates the watermark message embedding procedure, with the distinction that inverse rules are employed during the message extraction phase. In order to conduct semantic modeling more effectively, Munyer and Zhong (2023) utilized a pretrained Word2Vec model instead of WordNet to find alternatives. In particular, they converted the selected words into Word2Vec vectors and collected the n-nearest vectors to form replacement word set. Noticeably, they trained a binary classifier as watermark detector, using pretrained BERT model and transformer blocks as neural network components.
However, the aforementioned watermarking approaches, which depended on context-independents synonym substitution (WordNet & Word2Vec), tend to overlook the context of the target words when generating substitute candidates. This oversight may result in a failure to preserve the overall semantics of the sentence, consequently diminishing the quality of the text. In response to this issue, context-aware lexical substitution is introduced into text watermarking techniques. Yang et al. (2022) proposed a novel BERT-based infill model to generate lexical substitution candidates, taking the overall sentence’s meaning into account. The watermark detection algorithm mimics the watermark generation process, first locating words containing the embedded watermark message, then generating substitute candidates, and finally applying the inverse embedding rules to extract the watermark message. To simplify the watermark detection process, Yang et al. (2023) developed a watermarking scheme by first computing a random binary encoding for each word, then replacing the words representing bit-0 with context-based synonyms that represent bit-1. Since the encoding computed for non-watermarked text adhere to a Bernoulli distribution and the distribution is altered during the watermarking process, statistical tests can be employed directly to detect the presence of the watermark. To further improve the robustness of watermark algorithms against watermark removal attacks, a study by Yoo et al. (2023b) fine-tuned BERT-based infill model with keyword-preserving and syntactically-invariant corruptions, which achieved a state-of-the-art robustness compared to previous approaches.
Lexical-based watermarking approaches embed watermarks by substituting synonyms in the text. However, the scope for semantically non-altering synonym replacements that do not affect the text structure is likely limited. Consequently, the capacity for the watermarks in lexical-based approaches is restricted, often necessitating a trade-off with text quality.

### v4

@@ Abstract.
Text watermarking algorithms are crucial for protecting the copyright of textual content. Historically, their capabilities and application scenarios were limited. However, recent advancements in large language models (LLMs) have revolutionized these techniques. LLMs not only enhance text watermarking algorithms with their advanced text understanding and generation abilities but also create a need for employing these algorithms to protect their own copyrights. This paper presents a comprehensive survey of text watermarking technology in the LLM era, aiming to further its development and application.
Text Watermark, Large Language Models, Copyright Protection
$c$s: Security and privacy Digital rights managementccs: Computing methodologies Natural language processing
@@ 1. Introduction
Text watermarking involves embedding unique, imperceptible identifiers (watermarks) into textual content. These watermarks are designed to be robust yet inconspicuous, ensuring that the integrity and ownership of the content are preserved without affecting its readability or meaning. Historically, text watermarking has played a crucial role in various domains, from copyright protection and document authentication to preventing plagiarism and unauthorized content distribution (Kamaruddin et al., 2018). With the advancement of Large Language Models (LLMs), both the techniques and application scenarios of text watermarking have seen significant development. As shown in Figure 1(a), this primarily includes the construction of enhanced text watermarking algorithms using LLMs, the application of existing text watermarking algorithms to LLMs, and the exploration of text watermarking algorithms that are more closely integrated with LLMs. The flourishing development of LLMs has propelled a thriving research landscape within the realm of text watermarking, as depicted in Figure 1(b). Especially with the advent of ChatGPT, text watermarking has notably surged into a research fervor. This paper surveys the interplay between LLMs and text watermarking, highlighting their synergistic potential in the contemporary landscape of language models.
In the following content, we will separately discuss why text watermarking benefits the application of LLMs (section 1.1), why utilizing LLMs can lead to the development of superior text watermarking algorithms (section 1.2), and the contributions of this survey (section 1.3).
@@ 1.1. Why is Text Watermarking Beneficial for LLMs?
However, the utilization of Large Language Models (LLMs) introduces several challenges. Primarily, their capacity for rapid, high-quality text generation may expedite the dissemination of misinformation (Megías et al., 2021). Additionally, there are critical intellectual property concerns involving LLMs, encompassing both the copyright of training datasets (Tang et al., 2023) and the establishment of rights to impede knowledge extraction from LLMs (Zhao et al., 2023b). Implementing effective tagging and detection methods for LLM-generated text would substantially alleviate these problems. Text watermarking presents as a viable approach to these challenges. It involves embedding a discrete, identifiable marker within LLM-produced text, enabling content tracking and origin attribution.
@@ 1.2. Why are LLMs Beneficial for Text Watermarking?
A key challenge in text watermarking is to embed watermarks without distorting the original text’s meaning or readability. Traditional methods often fail to modify text without altering its semantics (Atallah et al., 2001; Topkara et al., 2006a; Meral et al., 2009). The necessity for algorithms to comprehend and control text semantics contributes to this difficulty. However, Large Language Models (LLMs) significantly alter this landscape. Due to their advanced grasp of language semantics and context, LLMs facilitate sophisticated watermarking approaches that embed watermarks with minimal impact on the text’s inherent meaning (Abdelnabi and Fritz, 2021; Zhang et al., 2023a). This integration results in more effective and subtle watermarking techniques, preserving the text’s original intent while embedding essential watermark features.
@@ 1.3. Why a Survey for Text Watermarking in the Era of LLMs?
Text watermarking technology and large language models can effectively enhance each other. For instance, text generated by LLMs can be watermarked using text watermarking algorithms (Brassil et al., 1995; Por et al., 2012; Rizzo et al., 2016; Munyer and Zhong, 2023; Yang et al., 2022, 2023; Yoo et al., 2023b), or LLMs themselves can be utilized to embed watermarks in texts (Abdelnabi and Fritz, 2021; Zhang et al., 2023a). Additionally, watermark algorithms can be directly incorporated during the text generation process of LLMs (Kirchenbauer et al., 2023a; Zhao et al., 2023a; Liu et al., 2023c, b; Ren et al., 2023; Wu et al., 2023). However, comprehensive studies exploring text watermarking in the era of LLMs are lacking. Existing surveys predominantly focus on watermarking techniques developed before the advent of LLMs (Alkawaz et al., 2016; Kamaruddin et al., 2018).
In this study, we present the first comprehensive survey of text watermarking algorithms in the context of large language models. This survey encompasses a detailed definition of text watermarking algorithms and explores the interrelations among various methods. Recognizing the complexity and diversity of text watermarking technology, we examine the evaluation metrics for these algorithms, focusing on success rate, robustness, impact on text quality, and unforgeability. Moreover, we discuss current applications of text watermarking, such as copyright protection, fake news detection, and academic integrity. This survey aims to furnish researchers with an in-depth understanding of text watermarking algorithms, facilitating comparisons and contrasts among different methods. It also guides users interested in applying text watermarking technology by helping them choose suitable algorithms under different application scenarios.
Organization of this survey. This survey is structured as follows: Section 2 introduces text watermarking definitions and key algorithm properties. Section 3 and Section 4 address two primary text watermarking categories: for existing text and for LLM-generated text. Section 5 discusses evaluation metrics for these algorithms, including success rate, impact on text quality, robustness, and unforgeability. Section 6 explores application scenarios, namely copyright protection, academic integrity, and fake news detection. Section 7 examines ongoing challenges and potential future research avenues in text watermarking. The survey concludes in Section 8.
@@ 2.1. Text Watermarking Algorithms
This watermarked text $\mathbf{t}$ is either an altered form of the original text $\mathbf{x}$ or a newly generated text in response to $\mathbf{x}$, particularly in contexts like prompts for LLMs. The watermark message, denoted as $w$, can be a zero-bit watermark, signifying merely its presence or absence, or a multi-bit watermark, embedding detailed, customized information. The phrase ‘watermark payload’ will henceforth denote the information volume conveyed by $w$.
@@ 2.2. Key Characteristics of Text Watermarking Algorithms
To enhance the understanding of objectives in text watermarking algorithm design, this section highlights the critical characteristics these algorithms should have. These primarily include a high success rate of watermark detection, minimal impact on text quality, robustness of the detector against text modifications, and unforgeability.
@@
Low Impact on Text Quality. Let $\mathcal{A}(\mathbf{x},\text{None})$ represent the text generated without watermark. When $\mathbf{x}$ is the target text, the output remains $\mathbf{x}$. In the context of a prompt for a LLM, it denotes the LLM’s output without watermark insertion. An effective watermark algorithm minimally affects text quality, satisfying the condition:
@@
where $\mathcal{R}$ is a function evaluating text quality from multiple perspectives, as will be discussed in section 5. $\delta$ represents a threshold. If the difference in the evaluated scores of two texts is less than this threshold, they are considered to be of similar quality.
Robustness to watermark removal attack. We use an operation $\mathcal{U}$ to denote the watermark removal operations, which will be detailed in section 5. If a watermarking algorithm is robust against watermark removal attacks, it should satisfy the following conditions:
@@
Unforgeability. Unforgeability refers to the difficulty for a third party to counterfeit text watermarks. Typically, if the watermark’s generator $\mathcal{A}$ is acquired by an attacker, the watermark can certainly be forged. Consequently, unforgeability primarily hinges on whether an attacker, lacking access to $\mathcal{A}$, can still counterfeit the watermark. This divides into two scenarios: the private detection scenario, where the attacker is devoid of the detector $\mathcal{D}$ and limited to detecting watermarks, and the public detection scenario, where the attacker possesses $\mathcal{D}$.
@@ 2.3. Taxonomy of Text Watermarking Algorithms
Figure 2 categorizes text watermarking algorithms into two primary types. The first, Watermarking for Existing Text, embeds watermarks into pre-existing texts, as elaborated in Section 3. This technique typically utilizes semantically invariant transformations for watermark integration.
The second type, Watermarking for Large Language Models, involves modifying LLMs, further detailed in Section 4. This method either embeds specific features into the training dataset or alters the LLMs’ text generation process, producing watermarked text from the input prompt.
@@ 3. Watermarking for Existing Text
Figure 3. This figure offers a concise overview of various text watermarking techniques. It categorizes watermarking into two main types: for Existing Text and for Large Language Models (LLMs). The former (section 3) embeds watermarks in existing text using methods like format-based (section 3.1), e.g., white-space substitution; lexical-based (section 3.2), e.g., synonym substitution; syntactic-based (section 3.3), e.g., passivization; and generation-based (section 3.4), employing pretrained language models. The latter (section 4) focuses on embedding watermarks in LLMs, ensuring their presence in generated text. This can be achieved during training (section 4.1), logits generation (section 4.2), or token sampling (section 4.3).
@@ 3.1. Format-based Watermarking
Format-based watermarking, drawing inspiration from image watermarking (Begum and Uddin, 2020), alters text format rather than content to embed watermarks. For example, Brassil et al. (1995) introduced line-shift coding and word shift-coding by vertically and horizontally adjusting text lines and words, respectively. The detection process measures distances between text line profiles or word column profiles to identify shifts. However, this method is confined to image-formatted text and does not return a text string with an embedded watermark. To address this, Unicode codepoint insertion/replacement methods have emerged. Por et al. (2012) developed UniSpach, inserting Unicode space characters in various text spacings. Rizzo et al. (2016) presented a unicode homoglyph substitution method, exploiting visually similar but differently coded text symbols (e.g., U+0043 and U+216d for ‘C’, U+004c and U+216c for ‘L’). Following this, a family of simple watermarks named EasyMark (Sato et al., 2023) is proposed recently, composed of three different methods: WhiteMark, VariantMark and PrintMark. Specifically, WhiteMark replaces a whitespace (U+0020) with another codepoint of a whitespace (e.g. U+2004). VariantMark leverages variation selectors of Unicode to embed hard-to-perceive format into CJK texts. PrintMark, coping with printed texts, uses ligature or whitespaces with slightly different lengths to embed watermark messages. Correspondingly, the watermark detection process involves searching for and counting the certain codepoints that have been inserted within the text. As these watermarking methods relied on the richness of Unicode encoding, their watermark payload is often quite large.
Despite the simplicity and effectiveness of format-based watermarking methods in embedding substantial payloads without altering textual content, their format modifications can be conspicuous under certain conditions, as Por et al. (2012) noted with the DASH attack. Consequently, these modified formats are vulnerable to removal via canonicalization (Boucher et al., 2022), involving techniques such as resetting line spacing and replacing specific codepoints throughout the text. Moreover, these detectable formats may be exploited for watermark forgery, compromising detection efficacy.
@@ 3.2. Lexical-based Watermarking
Format-based watermarking approaches, which only modify text’s superficial format, are prone to be spotted, making them easily removable through reformatting. This highlights the need for investigating deeper watermark embedding methods in text. A number of studies have adopted word-level modifications, where selected words are replaced with their synonyms without altering the sentence’s syntactic structure (Topkara et al., 2006b; Fellbaum, 1998; Munyer and Zhong, 2023; Yang et al., 2022, 2023; Yoo et al., 2023b). These are known as lexical-based watermarking approaches. Topkara et al. (2006b) introduced a synonym substitution method, employing WordNet (Fellbaum, 1998) as the synonym source. The watermark detection replicates the embedding process, applying inverse rules for message extraction. Munyer and Zhong (2023) enhanced semantic modeling by using a pretrained Word2Vec model, converting selected words into vectors and identifying n-nearest vectors as replacement candidates. They employed a binary classifier with a pretrained BERT model and transformer blocks for watermark detection.
The aforementioned watermarking methods relying on context-independent synonym substitution (WordNet & Word2Vec) often neglect the context of target words, potentially compromising sentence semantics and text quality. To address this, context-aware lexical substitution has been incorporated into text watermarking. Yang et al. (2022) introduced a BERT-based infill model for generating contextually appropriate lexical substitutions. The watermark detection algorithm parallels the generation process, identifying watermark-bearing words, generating substitutes, and applying inverse rules for message extraction. Yang et al. (2023) streamlined watermark detection by encoding each word with a random binary value and substituting bit-0 words with context-based synonyms representing bit-1. As non-watermarked text adheres to a Bernoulli distribution, altered during watermarking, statistical tests can effectively detect watermarks. Yoo et al. (2023b) enhanced robustness against removal attacks by fine-tuning a BERT-based infill model with keyword-preserving and syntactically invariant corruptions, achieving superior robustness compared to earlier methods.